OpenStudy (anonymous):

***I will give you a medal (: *** Write the equation in logarithmic form. 125^4/3=625 @kropot72

OpenStudy (anonymous):

a^b=c --> \[\log_{a} c=b\] So, \[\log_{125} 625=\dfrac {4} {3}\]

OpenStudy (anonymous):

is that the final answer ?

OpenStudy (mertsj):

Remember: the log is the exponent. So we can say that log = 4/3

OpenStudy (anonymous):

That's logarithmic form xD

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

Clearly the base is 125 so we can write \[\log_{125}=\frac{4}{3} \]

OpenStudy (mertsj):

That leaves only one place for the 625: \[\log_{125}625=\frac{4}{3} \]